A 75.0 kg hockey player moving at a speed of + 10.0 m / s collides with a second hockey player, who is motionless. After the collision, the two skaters move together at a speed of + 4.50 m / s. During the collision, what is the impulse received by the 75.0 kg player?


1
Expert's answer
2021-01-10T18:29:09-0500

By the definition, the impulse received by the 75.0 kg player equals the change in momentum of the 75.0 kg player:


"J=m\Delta v=m(v_f-v_i),""J=75.0\ kg\cdot(4.5\ \dfrac{m}{s}-10.0\ \dfrac{m}{s})=-412.5\ N\cdot s."

The sign minus means that the impulse received by the 75.0 kg player directed in the opposite direction to the motion of the 75.0 kg hockey player.

Answer:

"J=412.5\ N\cdot s," in the opposite direction to the motion of the 75.0 kg hockey player.